Given a straight wire with a 1A current flowing on it. Calculate the magnetic field magnitude at a distance of 3mm from the

wire.


1
Expert's answer
2021-06-11T11:28:43-0400

magnetic field;

"B=\frac{\mu_o I}{2\pi d}"

"\mu_o=4\pi \times 10^{-7}, I =1A, d=3mm=3\times 10^{-3}m"

"B=\frac{4\pi \times 10^{-7}\times 1}{2\pi \times3\times 10^{-3}}=6.67\times 10^{-5}T"